Amy paid $63.36 for a pair of running shoes during a 15% off sale. What was the regular price?
Train A and B are traveling in the same direction on parallel tracks. Tran A is traveling at 100 miles per hour and train B is traveling at 125 miles per hour. Train A passes a station at 1:25p.m If train B passes the same station at 1:37p.m., at what time will train B catch up to train A?

Answer by checkley77(12569)

63.36=(1-.15)x
63.36=.85x
x=63.36/.85
x=$74.54 was the original price.
======================================
D=RT
SEEING AS THE DISTANCES ARE EQUAL THEN:
100T=125(T-.2)
100T=125T-25
100T-125T=-25
-25T=-25
T=-25/-25
T=1 HOUR THEY WILL MEET.